### v1

# Efficient Generative Model Training via Embedded Representation Warmup

Abstract
Diffusion models excel at generating high-dimensional data but fall short in training efficiency and representation quality compared to self-supervised methods. We identify a key bottleneck: the underutilization of high-quality, semantically rich representations during training notably slows down convergence. Our systematic analysis reveals a critical representation processing region—primarily in the early layers—where semantic and structural pattern learning takes place before generation can occur. To address this, we propose Embedded Representation Warmup (ERW), a plug-and-play framework where in the first stage we get the ERW module serves as a warmup that initializes the early layers of the diffusion model with high-quality, pretrained representations. This warmup minimizes the burden of learning representations from scratch, thereby accelerating convergence and boosting performance. Our theoretical analysis demonstrates that ERW’s efficacy depends on its precise integration into specific neural network layers—termed the representation processing region—where the model primarily processes and transforms feature representations for later generation. We further establish that ERW not only accelerates training convergence but also enhances representation quality: empirically, our method achieves a 40$\times$ acceleration in training speed compared to REPA [54], the current state-of-the-art methods. Code is available at https://github.com/LINs-lab/ERW.
Figure 1: Embedded Representation Warmup (ERW). Throughout the training process, we demonstrate that incorporating representations at the early stages is highly beneficial. To this end, we propose a representation warmup stage that employs a representation alignment loss to integrate representations from models such as Dinov2  [38] into the ERW. This initialized representation region is subsequently embedded into the diffusion model pipeline, providing a strong starting point for training. Our ERW method thus enhances both efficiency and effectiveness, leading to faster convergence and superior performance compared to the REPA [54] method, thereby establishing a new state-of-the-art.
1 Introduction
“All roads lead to Rome, but it is not as good as being born in Rome.”
Deep generative models, particularly diffusion models [16, 49], have made remarkable strides in producing high-fidelity, realistic images. These models have demonstrated impressive performance in tasks such as unconditional image generation [8] and text-to-image synthesis [43, 45], largely due to their capacity to model complex data distributions through iterative denoising processes. By refining noisy inputs over multiple stages, diffusion models effectively capture intricate features that yield visually coherent outputs.
Despite these achievements, diffusion models often lag behind self-supervised learning methods in both training efficiency and the quality of learned representations. This gap is notable because efficient training and robust representation learning are pivotal for scaling generative models to real-world applications. While diffusion models implicitly learn semantic features during the denoising process [53, 52], these representations typically lack the depth and versatility seen in self-supervised approaches [3, 38]. Recent studies [29, 54] have investigated techniques to enhance internal representations in diffusion models, aiming to improve both the expressiveness of the generated data and the training efficiency. Moreover, Kadkhodaie et al. [20] highlight a critical bottleneck: the tension between memorizing semantic information and generalizing to realistic data distributions.
These observations lead to a pivotal question:
$\mathbb{Q}$: Can we first learn high-quality external representations from a well-established representation model and then integrate their representation into diffusion models, rather than training both representation and generation from scratch?
Self-supervised learning approaches, such as contrastive learning [4] and masked autoencoders [14], as well as more recent methods like Dinov2 [38], have shown strong capabilities in extracting rich semantic representations from unlabeled data. However, adapting these external representations to diffusion models poses inherent challenges. Primarily, diffusion models operate on progressively denoised inputs, whereas self-supervised encoders are trained on clean data. In addition, the architectures of self-supervised encoders may differ significantly from typical diffusion model backbones, further complicating direct integration.
Our approach. We observe that diffusion model training can be effectively conceptualized as a two-phase process: first learning semantic representations, and then leveraging those representations for generative purposes. Motivated by this viewpoint, we propose Embedded Representation Warmup (ERW), which explicitly separates training into two phases. In the warmup phase, we initialize the early-layer parameters of the diffusion model with weights from a pretrained self-supervised backbone (e.g., Dinov2). Since early layers predominantly capture semantic features, this partial initialization effectively injects strong semantic priors into the diffusion model. Our ablation studies confirm that restricting representation learning to early layers, rather than extending it to the entire architecture, leads to both better performance and higher training efficiency.
Furthermore, we decompose the diffusion model’s workflow into two distinct circuits: ① the Latent-to-Representation (L2R) circuit, mapping raw inputs into a semantic embedding space, and ② the Representation-to-Generation (R2G) circuit, synthesizing final outputs from these embeddings. Building on this insight, ERW uniformly initializes the model’s early layers to embed robust semantic features, forming a solid foundation for subsequent generative training.
Validation. Extensive experiments confirm that providing high-quality representations from the outset via ERW substantially accelerates convergence. We demonstrate a significant training speedup (up to $\mathbf{40\times}$ compared to REPA) while maintaining state-of-the-art generative quality. Notably, ERW enables SiT-XL to reach an FID of 6.4 on ImageNet-1k without classifier-free guidance in under 4 hours on 8$\times$×H800 GPUs. Since the warmup phase accounts for only a fraction of the overall training cost, reusing warmup-optimized checkpoints further enhances training efficiency across multiple tasks.
Contributions:
  1. (a)
    We identify and demonstrate that diffusion model training can be effectively decomposed into two distinct phases: a representation learning phase and a generation phase, with the representation processing mainly occurring in early layers (see  Section 3).
  2. (b)
    We introduce Embedded Representation Warmup, a two-phase training framework that first initializes the representation layers using pretrained self-supervised models before proceeding to the main generation-focused training (see  Section 4).
  3. (c)
    We empirically validate our approach on and demonstrating significant improvements in training efficiency and generation quality (see  Section 5).
2 Related Work
Our work builds upon and extends several areas of research in diffusion models, representation learning, and the integration of external representations for generative modeling.
Integrating pretrained representations into diffusion models.
Recent work exploits pretrained encoders to enhance diffusion-based generation. Würstchen [41] adopts a two-stage design where an initial network learns a semantic map, guiding a second diffusion model for high-resolution output. RCG [32] further underscores the utility of compact latent codes, generated by a first diffusion model and consumed by a second, boosting generative performance. REPA [54] applies regularization throughout the diffusion training process to align model representations with pretrained visual encoders. In contrast, our approach builds on these ideas by decoupling representation learning from the denoising process. We explicitly separates the training into two phases first learning representation then learning generation. This design not only alleviates the optimization challenges associated with joint learning but also leads to a significant acceleration in training convergence.
Leveraging self-supervised learning for generative modeling.
Self-supervised encoders pretrained via contrastive learning [4] or masked reconstruction [14] have made significant inroads in representation quality. Recent developments integrate such encoders into generative pipelines, for instance using them as discriminators in GAN-based frameworks [47, 26] or to guide diffusion distillation [30].
Enhancing representation learning within diffusion models.
Numerous studies highlight the discriminative potential of diffusion representations [53, 52], revealing that intermediate layers encode semantic features. Efforts like Repfusion [53] and DreamTeacher [30] distill these features to smaller networks for downstream tasks. Inspired by these findings, we identify a distinct representation processing region where crucial representation learning occurs before generation can take place. By initializing this region with pretrained weights during a dedicated warmup phase, we circumvent the costly need to learn complex features from scratch in the main training phase, leading to faster convergence and improved fidelity.
Towards unified and efficient diffusion training.
Beyond two-stage pipelines [41, 32], alternative methods fuse self-supervision with the diffusion objective, such as MaskDiT [55] or MAGE [31], yet often require careful objective balancing. Our plug-and-play method instead introduces a clear two-phase approach: a representation-focused warmup phase that initializes early diffusion layers with mature pretrained weights, followed by a standard diffusion training phase that benefits from this strong representational foundation.
Figure 2: Illustration of the Circuits in Hypothesis. From left to right, we first apply a VAE encoder (the Pixel-to-Latent or (P2L) stage) to map high-dimensional inputs to a compressed latent space. We then perform latent diffusion on these codes, dividing the backbone into two subregions: the Latent-to-Representation (L2R) region that capture and refine semantic features, and the Representation-to-Generation (R2G) region that decode the learned representation into final outputs.
3 Hypothesis of Latent Diffusion Models
In this section, we propose a three-stage diffusion circuit—Pixel-to-Latent (P2L), Latent-to-Representation (L2R), and Representation-to-Generation (R2G)—as a structured framework for more effective and interpretable generative modeling. Our guiding hypothesis is that conventional diffusion methods entangle representation learning and generative decoding within the denoising process, burdening the network with two inherently distinct tasks. By explicitly isolating the semantic processing (L2R) from the generative inversion (R2G), our approach promotes modularity and simplifies the overall optimization landscape. This theoretical perspective also clarifies how different regions of the diffusion backbone specialize in semantic vs. generative aspects, offering a fresh viewpoint on the internal mechanics of diffusion models.
3.1 Preliminaries
Latent diffusion models.
While classic diffusion models such as DDPM [16] adopt a discrete-time denoising process, flow-based methods [33, 1, 48] explore diffusion in a continuous-time setting. In particular, Scalable Interpolant Transformers (SiT) [10, 33, 34] offer a unifying framework for training diffusion models on a continuous-time stochastic interpolant. Below, we describe how SiT can be leveraged to learn powerful latent diffusion models.
Forward process via stochastic interpolants.
Consider a data sample $\mathbf{x}\sim p(\mathbf{x})$ (e.g., an image) and let the encoder $\mathcal{H}_{\bm{\theta}}(\mathbf{x})$ map it to its latent representation denoted as $\mathbf{z}_{\ast}\in\mathcal{Z}$. Given standard Gaussian noise $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, SiT defines a forward process in the latent space, parameterized by continuous time $t\in[0,1]$:
$\mathbf{z}_{t}=\alpha_{t}\,\mathbf{z}_{\ast}\,+\,\sigma_{t}\,\bm{\epsilon}\,,$ (1)
where $\alpha_{t}$ and $\sigma_{t}$ are deterministic, differentiable functions satisfying the boundary conditions:
$(\alpha_{0},\sigma_{0})=(1,0)\quad\text{and}\quad(\alpha_{1},\sigma_{1})=(0,1)\,.$ (2)
This construction implies that $\mathbf{z}_{0}=\mathbf{z}_{\ast}$ (the clean latent representation) and $\mathbf{z}_{1}=\bm{\epsilon}$ (pure noise). Under mild conditions [1], the sequence $\{\mathbf{z}_{t}\}$ forms a stochastic interpolant that smoothly transitions between data and noise in the latent space.
Velocity-based learning.
To train a diffusion model in this continuous-time framework, SiT employs a velocity formulation. Differentiating $\mathbf{z}_{t}$ with respect to $t$ yields:
$\dot{\mathbf{z}}_{t}=\dot{\alpha}_{t}\,\mathbf{z}_{\ast}\,+\,\dot{\sigma}_{t}\,\bm{\epsilon}\,.$ (3)
Conditioning on $\mathbf{z}_{t}$, we can rewrite the derivative as a velocity field:
$\dot{\mathbf{z}}_{t}=\mathbf{v}(\mathbf{z}_{t},t)\,,$ (4)
where $\mathbf{v}(\mathbf{z}_{t},t)$ is defined as the conditional expectation of $\dot{\mathbf{z}}_{t}$ given $\mathbf{z}_{t}$. A neural network $\mathbf{v}_{\bm{\theta}}(\mathbf{z},t)$ is then trained to approximate $\mathbf{v}(\mathbf{z},t)$ by minimizing:
$\small\textstyle\mathcal{L}_{\mathrm{velocity}}(\bm{\theta})\;=\;{\mathbb{E}}_{\mathbf{z}_{\ast},\bm{\epsilon},t}\left[\left\lVert\mathbf{v}_{\bm{\theta}}(\mathbf{z}_{t},t)-\Bigl{(}\dot{\alpha}_{t}\,\mathbf{z}_{\ast}+\dot{\sigma}_{t}\,\bm{\epsilon}\Bigr{)}\right\rVert^{2}\right]\,.$ (5)
Successfully learning $\mathbf{v}_{\bm{\theta}}(\mathbf{z},t)$ enables one to integrate the reverse-time ordinary differential equation (ODE) [49], thereby mapping noise samples back to coherent latent representations.
3.2 Three-Stage Diffusion Circuit
Recent studies indicate that diffusion models jointly perform both representation learning and generative decoding during the denoising procedure [54, 52]. Notably, every layer in the network contributes to feature extraction and generative tasks to varying degrees. To make this dual functionality clearer, we propose decomposing the diffusion process into three distinct stages: Pixel-to-Latent (P2L), Latent-to-Representation (L2R), and Representation-to-Generation (R2G), as illustrated in Figure 2. Formally, we posit that the diffusion sampling procedure can be written as:
$\mathbf{z}$ $=\mathcal{H}_{\bm{\theta}}(\mathbf{x})\,,$ (Pixel to Latent (P2L))
$\mathbf{r}$ $=\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z})\,,$ (Latent to Representation (L2R))
$\mathbf{z}^{\prime}$ $=\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}(\mathbf{r})\,,$ (Representation to Generation (R2G))
Here, the Variational Autoencoder (VAE) encoder $\mathcal{H}_{\bm{\theta}}$ compresses high-dimensional pixel data $\mathbf{x}$ into a lower-dimensional latent $\mathbf{z}$. The diffusion model components $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ and $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ share parameters in diffusion model backbone but serve distinct functions: $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ extracts semantic representations from noisy latent codes (L2R), while $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ transforms these representations into refined latent outputs (R2G). During inference, a VAE decoder $\mathcal{D}_{\bm{\theta}}$ would convert these refined latents back to pixel space. This decomposition into distinct processing stages provides a more transparent view of how latent diffusion models function and support the idea from latent space compression in Saharia et al. [45] and the diffusion model training via represention alignment in REPA Yu et al. [54].
Loss Function Decomposition.
The diffusion model’s objective is to learn the score function $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})$, which represents the gradient of the log probability density of the data distribution. Our key theoretical insight is that this score function can be orthogonally decomposed into representation and generation components (see details in Appendix A):
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})=\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})}_{\mathrm{Representation}}+\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})}_{\mathrm{Generation}}\,.$ (6)
This decomposition permits us to partition the overall diffusion objective into two nearly orthogonal terms, focusing respectively on (i) extracting meaningful features from corrupted inputs and (ii) translating those features into coherent outputs. Such a structure naturally aligns with the three-stage diffusion circuit we will soon discuss, clarifying the distinct roles played by different network regions.
Stage I: Pixel-to-Latent (P2L).
Before performing the denoising process in the high-dimensional pixel domain—where noise may obscure semantic cues—many contemporary methods [45, 16, 8] compress images into a more tractable latent space:
$\mathbf{z}\;=\;\mathcal{H}_{\bm{\theta}}(\mathbf{x})\,,$ (7)
where $(\mathcal{H}_{\bm{\theta}},\mathcal{D}_{\bm{\theta}})$ typically refers to a variational autoencoder or a related autoencoding architecture. This P2L stage reduces computational complexity and filters out low-level details, thus preserving more essential semantic information. From the perspective of the decomposed loss, P2L transforms the high-dimensional denoising problem into a lower-dimensional one where representation components (capturing semantic concepts) and reconstruction components (handling fine details) become more clearly separable, facilitating favorable conditions for separating the training stages.
Figure 3: Selected Samples on ImageNet $256\times 256$. Images generated by the SiT-XL/2 + REPA + ERW model using classifier-free guidance (CFG) with a scale of $w=2.2$ under 40 epochs.
Stage II: Latent-to-Representation (L2R).
Given a noisy latent $\mathbf{z}_{t}$ from Eq. (7), the model initially extracts a semantic representation $\mathbf{r}_{t}$ using the mapping $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$.
$\mathbf{r}_{t}\;=\;\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t})\,.$ (8)
This step corresponds to estimating $\nabla\log p(\mathbf{r}_{t})$ within the velocity loss decomposition (where we separate representation learning from generation, as detailed in  Appendix A), sometimes termed the Representation Score [49]. Intuitively, the model should discern salient patterns (e.g., object shapes, style characteristics, or conditioning signals) before denoising. Under the sufficient statistic assumption (see Assumption 1), the representation $\mathbf{r}_{t}$ effectively captures the essential information from the latent $\mathbf{z}_{t}$, allowing us to formulate the representation learning process as an independent objective:
$\mathcal{L}_{\mathrm{L2R}}\;=\;\mathbb{E}_{t,\mathbf{z}_{t}}\bigl{[}\|\,\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t})\!-\!\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})\,\|^{2}\bigr{]}\,,$ (9)
where $\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}$ represents the parameterized representation extraction module of the network. By explicitly decoupling the objective for semantic feature extraction from that of generative refinement, the model is guided to learn robust high-level representations and ensures that the early layers focus on capturing essential semantic features.
Stage III: Representation-to-Generation (R2G).
In the final phase of each reverse diffusion update in (3), known as the R2G stage, the extracted semantic representation is transformed into an updated latent with reduced noise:
$\mathbf{z}_{t-\Delta t}=\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}(\mathbf{r}_{t},t)\,.$ (10)
This output serves the same purpose as the $\mathbf{z}^{\prime}$ term introduced in our framework’s third Eq. (3), but is specifically defined for the discrete time step $t-\Delta t$ in the continuous-time diffusion process. In the decomposed view, this step aligns with the Generation Score term, $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})$, detailed in  Appendix A:
$\mathcal{L}_{\mathrm{R2G}}\!=\!\mathbb{E}_{t,\mathbf{z}_{t}}\Big{[}\left\lVert\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}(\mathbf{r}_{t},t)\!-\!\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\!\mid\!\mathbf{r}_{t})\right\rVert^{2}\Big{]}\,,$ (11)
where $\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}$ represents the generation module’s score estimator, transforming representations back into the latent space for denoising. Injecting the semantic representation $\mathbf{r}_{t}$ into a cleaner latent $\mathbf{z}_{t-\Delta t}$—which is significantly less noisy than $\mathbf{z}_{t}$—ensures that abstract semantic features are effectively transformed into the precise latent elements required for content generation. Meanwhile, the cross term between L2R and R2G (also elaborated in  Appendix A) remains small under the assumption that these two gradients are sufficiently orthogonal, mitigating destructive interference.
Two-Stage Sampling Process: Representation Extraction Precedes Generation.
Within the continuous-time framework, after mapping pixel data to the latent space through P2L, each infinitesimally small time step during the reverse SDE update can be interpreted as a two-stage process:
$\mathbf{r}_{t}=\mathcal{R}_{\bm{\theta}_{\mathrm{L2R}}}(\mathbf{z}_{t})\quad\longrightarrow\quad\mathbf{z}_{t-\Delta t}=\mathcal{G}_{\bm{\theta}_{\mathrm{R2G}}}(\mathbf{r}_{t})$ (12)
Hence, every time step naturally splits into (i) L2R for refining the representation and (ii) R2G for synthesizing an updated latent. This loop neatly implements the principle of “first representation, then generation”. Empirically, prior work [54, 52] confirms that early layers of the diffusion model predominantly focus on representation extraction, whereas later layers emphasize generative refinement. Consequently, the staged design mirrors the reverse-time diffusion trajectory, concluding in a final latent $\mathbf{z}_{0}$ that is decoded via $\mathcal{D}_{\bm{\theta}}$ to yield the synthesized output $\mathbf{x}_{0}$.
4 Embedded Representation Warmup: A Two-Phase Training Framework
Diffusion models generate high-fidelity images [16, 8], yet their training is impeded by the simultaneous challenges of learning robust semantic features and performing generative denoising [49]. This coupling creates an inherently complex optimization landscape that demands prolonged training periods.
In this section, we present Embedded Representation Warmup (ERW), a framework that strategically decouples these challenges into two distinct phases. In Phase 1, we initialize the early layers of the diffusion model with high-quality semantic features from pretrained models; in Phase 2, we transition to standard diffusion training with a gradually diminishing representation alignment term, allowing the model to increasingly focus on generation. This separation addresses the inefficiency of joint optimization in conventional diffusion models and significantly enhances training efficiency.
4.1 Phase 1: Representation Warmup Stage
To alleviate the burden of learning semantic features from scratch, we begin with a dedicated warmup stage. During this phase, the model’s L2R circuit is initialized to align with semantically rich features extracted from a pretrained representation model (e.g., DINOv2, MAE, or CLIP).
Let $\mathcal{H}_{\mathbf{\theta}}(\mathbf{x})$ denote an encoder that maps an image $\mathbf{x}\in\mathcal{X}$ to its latent representation $\mathbf{z}\in\mathcal{Z}$, and let $\mathbf{f}_{\mathrm{rep}}:\mathcal{X}\to\mathcal{R}$ be a high-quality pretrained representation model. Our objective is to learn a mapping $\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}$ such that its output closely aligns with $\mathbf{f}_{\mathrm{rep}}(\mathbf{x})$.
Specifically, we define the alignment loss as:
$\mathcal{L}_{\mathrm{align}}=\mathbb{E}_{\mathbf{x}\sim p_{\mathrm{data}}(\mathbf{x})}\Bigl{[}\bigl{\|}\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}(\mathcal{H}_{\mathbf{\theta}}(\mathbf{x}))-\mathbf{f}_{\mathrm{rep}}(\mathbf{x})\bigr{\|}_{2}^{2}\Bigr{]}$ (13)
By minimizing Eq. (13), we ensure that $\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}$ produces features well-aligned with the pretrained representation model. This alignment mitigates the need for the diffusion model to learn complex semantic features purely from noisy inputs; instead, subsequent layers can operate on features that are already semantically meaningful.
4.2 Phase 2: Generative Training with Decaying Representation Guidance
After the warmup stage has effectively initialized the diffusion model with semantically rich features, we proceed with a joint objective that combines the standard diffusion loss with a gradually diminishing representation alignment term. Formally, the overall training loss is given by:
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{diffusion}}+\lambda(t)\cdot\mathcal{L}_{\mathrm{align}}$ (14)
Here, $\mathcal{L}_{\mathrm{diffusion}}$ denotes the conventional diffusion loss (e.g., the velocity prediction loss as defined in Eq. (5)), $\mathcal{L}_{\mathrm{align}}$ is the representation alignment loss, and $\lambda(t)$ is a time-dependent weight that modulates the impact of alignment during training.
In this phase, the alignment loss $\mathcal{L}_{\mathrm{align}}$ encourages the diffusion model’s hidden representations, after being projected by a trainable function $\mathcal{T}_{\mathbf{\theta}}$, to maintain alignment with the high-quality semantic features of the external encoder:
$\mathcal{L}_{\mathrm{align}}=\mathbb{E}_{t,\mathbf{x}}\Biggl{[}\Bigl{\|}\mathcal{T}_{\mathbf{\theta}}\Bigl{(}\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}(\mathcal{H}_{\mathbf{\theta}}(\mathbf{x}))\Bigr{)}-\mathbf{f}_{\mathrm{rep}}(\mathbf{x})\Bigr{\|}_{2}^{2}\Biggr{]}$ (15)
The time-dependent alignment weight follows an exponential decay schedule:
$\lambda(t)=c_{0}\exp\!\left(-\frac{t}{\tau}\right)$ (16)
where $c_{0}$ is the initial weight of the alignment loss and $\tau$ is a hyperparameter controlling the decay rate. This schedule provides strong semantic guidance during early training—when representation learning is most critical—and gradually reduces its influence, allowing the network to increasingly focus on optimizing generative capabilities.
4.3 Theoretical Analysis: Warmup Leads to Faster Convergence
Our warmup phase minimizes the misalignment between the model’s internal representation $\mathcal{R}_{\mathbf{\theta}_{\mathrm{L2R}}}(\mathcal{H}_{\mathbf{\theta}}(\mathbf{x}))$ and the target semantic features $\mathbf{f}_{\mathrm{rep}}(\mathbf{x})$. After warmup, this misalignment becomes negligible for almost all data samples $\mathbf{x}$.
With the residual error minimized, the overall training objective during the diffusion stage is effectively dominated by the generative loss $\mathcal{L}_{\mathrm{diffusion}}$. The alignment loss $\mathcal{L}_{\mathrm{align}}$, scaled by the diminishing weight $\lambda(t)$, contributes minimally to the total loss. This enables the model to focus on refining its generative performance rather than correcting representation discrepancies.
To formalize this intuition, consider the gradient of the total loss with respect to the model parameters $\mathbf{\theta}$:
$\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{total}}=\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{diffusion}}+\lambda(t)\cdot\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{align}}$ (17)
Since the alignment loss has been minimized during warmup, its gradient $\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{align}}$ is small. Moreover, the alignment weight $\lambda(t)$ decays over time (as specified in Eq. (16)), further reducing the impact of the second term. Consequently, parameter updates are predominantly driven by the diffusion loss:
$\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{total}}\approx\nabla_{\mathbf{\theta}}\mathcal{L}_{\mathrm{diffusion}}$ (18)
This separation means the model no longer needs to allocate capacity to learn semantic features from scratch during the main training phase. Instead, it can concentrate on perfecting its generative capabilities, leading to substantially faster convergence and improved sample quality.
5 Experiments
In this section, we provide a comprehensive evaluation of our proposed ERW approach. We begin by outlining experimental setups ( Section 5.1), including dataset and implementation details. Next, we present comparisons with state-of-the-art baselines to demonstrate the benefits of ERW in both FID and training speed ( Section 5.2). We then analyze the role of our warmup procedure in boosting training efficiency ( Section 5.3). Finally, we conduct ablation studies to examine the effects of various alignment strategies, architecture depths, and target representation models ( Section 5.4).
5.1 Setup
$\bullet$ Implementation Details. We adhere closely to the experimental setups described in DiT [40] and SiT [36], unless otherwise noted. Specifically, we utilize the ImageNet dataset [7], preprocessing each image to a resolution of $256\times 256$ pixels. Following the protocols of ADM [8], each image is encoded into a compressed latent vector $\mathbf{z}\in\mathbb{R}^{32\times 32\times 4}$ using the Stable Diffusion VAE [44]. For our model configurations, we employ the B/2, L/2, and XL/2 architectures as introduced in the SiT papers, which process inputs with a patch size of 2. To ensure a fair comparison with SiT models, we maintain a consistent batch size of 256 throughout training. Further experimental details, including hyperparameter settings and computational resources, are provided in Appendix C.
$\bullet$ Evaluation. We report Fréchet inception distance (FID; Heusel et al. 15), sFID [37], inception score (IS; Salimans et al. 46), precision (Pre.) and recall (Rec.) [27] using 50K samples. We also include CKNNA [19] as discussed in ablation studies. Detailed setups for evaluation metrics are provided in Appendix D.
Table 1: System-level comparison on ImageNet 256$\times$256 with CFG. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively. Results that include additional CFG scheduling are marked with an asterisk (*), where the guidance interval from [28] is applied for ERW.
[…]
Pixel diffusion
  00ADM-U 0400 03.94 6.14 186.7 0.82 0.52
  00VDM+++++ + 0560 02.40 - 225.3 - -
  00Simple diffusion 0800 02.77 - 211.8 - -
  00CDM 2160 04.88 - 158.7 - -
Latent diffusion, U-Net
  00LDM-4 0200 03.60 - 247.7 0.87 0.48
Latent diffusion, Transformer + U-Net hybrid
  00U-ViT-H/2 0240 02.29 5.68 263.9 0.82 0.57
  00DiffiT* - 01.73 - 276.5 0.80 0.62
  00MDTv2-XL/2* 1080 01.58 4.52 314.7 0.79 0.65
Latent diffusion, Transformer
  00MaskDiT 1600 02.28 5.67 276.6 0.80 0.61
  00SD-DiT 0480 03.23 - - - -
  00DiT-XL/2 1400 02.27 4.60 278.2 0.83 0.57
  00SiT-XL/2 1400 02.06 4.50 270.3 0.82 0.59
  000+ REPA 0200 01.96 4.49 264.0 0.82 0.60
  $0$+ ERW (ours) 0040 01.94 7.91 300.9 0.75 0.63
$\bullet$ Sampler and Alignment objective. Following SiT [36], we always use the SDE Euler-Maruyama sampler (for SDE with $w_{t}=\sigma_{t}$) and set the number of function evaluations (NFE) as 250 by default. We use Normalized Temperature-scaled Cross Entropy (NT-Xent) training objective for alignment.
$\bullet$ Baselines. We use several recent diffusion-based generation methods as baselines, each employing different inputs and network architectures. Specifically, we consider the following four types of approaches: (a) Pixel diffusion: ADM [8], VDM+++++ + [23], Simple diffusion [18], CDM [17], (b) Latent diffusion with U-Net: LDM [44], (c) Latent diffusion with transformer+U-Net hybrid models: U-ViT-H/2 [2], DiffiT [12], and MDTv2-XL/2 [11], and (d) Latent diffusion with transformers: MaskDiT [55], SD-DiT [56], DiT [40], and SiT [36]. Here, we refer to Transformer+U-Net hybrid models that contain skip connections, which are not originally used in pure transformer architecture. Detailed descriptions of each baseline method are provided in   Appendix E.
5.2 Comparison to Prior Methods
Table 2: FID comparisons with SiT-XL/2. In this table, we report the FID of ERW with SiT-XL/2 on ImageNet $256\times 256$ at various Training iterations. Here is only full training without warmup, because we load a well trained warmuped checkpoint. For comparison, we also present the performance of the state-of-the-art baseline REPA at similar iterations or comparable FID values. Note that $\downarrow$ indicates that lower values are preferred and all results reported are without classifier-free guidance.
[…]
SiT-XL/2 675M 7M 8.3 6.32 131.7 0.68 0.67
+REPA 675M 50K 52.3 31.24 24.3 0.45 0.53
+ERW (ours) 675M 50K 8.5 9.59 154.7 0.72 0.51
+REPA 675M 100K 19.4 6.06 67.4 0.64 0.61
+ERW (ours) 675M 100K 6.0 6.37 207.5 0.67 0.63
Figure 4: Comparison of Training Efficiency and Cost Analysis with Warmup and Full Training Stages. Left: Scatter plot depicting the relationship between the total training cost (in TFLOPs) and the FID score for various training strategies. Each point is annotated with a simplified label (e.g., “10K+90K”) representing the warmup and full training split, and the marker sizes are scaled based on a combination of the FID and FLOPs values to highlight the relative differences. Right: Bar chart comparing the computational costs of the warmup and full training stages for different strategies (all evaluated over 100K iterations). The chart shows the warmup cost, full training cost, and their corresponding total cost.
Table 1 summarizes our results on ImageNet $256\times 256$ under classifier-free guidance (CFG). Our ERW significantly boosts the convergence speed of SiT-XL/2, enabling strong FID scores at just 40 epochs. Importantly, we observe that ERW provides an alternative path to large-scale diffusion modeling by separating representational learning from generative training, thereby reducing the need for extended epochs.   Figure 3 illustrates generated samples on ImageNet, further confirming the high-quality outputs achieved by ERW.
5.3 Effect of ERW: Quantitative Gains and Training Efficiency
Performance without Classifier-Free Guidance.
We begin by examining how ERW influences SiT-XL/2’s Frechet Inception Distance (FID) when no classifier-free guidance (CFG) is used.
  1. (a)
    Efficient FID Improvements. In  Table 5.2, ERW consistently achieves competitive or superior FID values compared to baselines. For instance, ERW reaches an FID of 6.0 at 100k iterations, markedly outperforming the REPA method [54] within the same iteration budget.
  2. (b)
    Leveraging Pretrained Features. This gain highlights the advantage of injecting pretrained semantic priors via warmup, thereby accelerating convergence in the subsequent full diffusion training.
Summary. By incorporating a short warmup step, ERW enables the model to focus its capacity on generative decoding rather than learning semantic features from scratch, resulting in improved FID at fewer training steps.
Warmup Versus Full Training.
Next, we analyze how splitting the total training budget between warmup and full diffusion training impacts both generation quality and computational overhead ( Figure 4).
  1. (a)
    Training Splits. We experiment with various allocations: 10k warmup + 90k full training, 20k warmup + 80k full training, and so on, always keeping the total at 100k iterations.
  2. (b)
    Accelerated Convergence. As shown in  Figure 4, a 10k warmup already yields a notable FID improvement (down to 6.4) with only a slight increase in total FLOPs ( Figure 4). This underscores that early alignment with a strong pretrained encoder can significantly hasten the later diffusion training.
Summary. A dedicated, relatively brief warmup phase materializes substantial downstream speedups and quality gains, confirming that investing in warmup alignment is a cost-effective strategy.
Table 3: Analysis of ERW on ImageNet 256$\times$256. All models are SiT-B/2 trained for 50K iterations. All metrics except FID without classifier-free guidance. For Acc., we report linear probing results on the ImageNet validation set using the latent features aligned with the target representation. We fix $\lambda=0.5$ here. $\downarrow$ and $\uparrow$ indicate whether lower or higher values are better, respectively.
[…]
MoCov3-B 8 NT-Xent 61.1 7.6 22.38 0.42 0.58
MoCov3-L 8 NT-Xent 73.0 8.0 17.96 0.38 0.52
CLIP-L 8 NT-Xent 58.9 7.7 23.68 0.44 0.54
DINOv2-B 8 NT-Xent 55.6 7.8 25.45 0.44 0.56
DINOv2-L 8 NT-Xent 55.5 7.8 25.45 0.44 0.56
DINOv2-g 8 NT-Xent 59.4 7.6 25.53 0.44 0.56
Table 4: Analysis of ERW places influences in SiT-B/2. All models are based on SiT-B/2 and trained for 50K iterations under the batch size of 256 without using classifier-free guidance on ImageNet $256\times 256$. $\downarrow$ indicates lower values are better. Results empirically validate our hypothesis that placing ERW at the forefront of the architecture yields optimal performance.
[…]
SiT-B/2 + REPA [54] 78.2 11.71 17.1 0.33 0.48
DINOv2-B 0-8 NT-Xent 54.2 8.12 27.2 0.45 0.59
DINOv2-B 1-9 NT-Xent 69.1 13.0 18.7 0.37 0.51
DINOv2-B 2-10 NT-Xent 67.7 13.4 19.0 0.38 0.52
DINOv2-B 3-11 NT-Xent 67.5 11.8 19.5 0.38 0.52
DINOv2-B 4-11 NT-Xent 67.8 13.1 19.0 0.38 0.52
Table 5: Analysis of ERW depth, projection depth, and different dynamic or consistent projection loss $\lambda$ influences in SiT-XL/2. All models are based on SiT-XL/2 and trained for 50K iterations under a batch size of 256 without using classifier-free guidance on ImageNet $256\times 256$. The target representation model is DINOv2-B, and the objective is NT-Xent. $\downarrow$ indicates lower values are better. The results show that a projection depth of 14 and a projection loss $\lambda$ of 4.0 yield substantial improvements in both FID and sFID, indicating an optimal configuration for model performance.
[…]
SiT-XL/2 + REPA [54] 52.3 31.24 24.3 0.45 0.53
3 8 0.5 13.3 13.65 113.2 0.69 0.48
4 8 0.5 11.7 17.93 114.5 0.68 0.48
5 8 0.5 13.1 15.25 117.0 0.69 0.48
6 8 0.5 13.7 16.62 115.7 0.68 0.49
8 8 0.5 15.6 16.42 104.0 0.67 0.46
12 8 0.5 19.2 16.91 87.8 0.65 0.48
4 8 0.5 13.7 17.93 114.5 0.68 0.48
4 10 0.5 12.8 17.25 121.2 0.68 0.48
4 12 0.5 12.3 17.03 123.7 0.69 0.48
4 14 0.5 12.4 16.97 124.4 0.68 0.48
4 16 0.5 12.6 16.79 124.2 0.67 0.49
4 14 0.1 21.6 19.58 74.9 0.62 0.49
4 14 0.5 12.4 16.97 124.3 0.68 0.48
4 14 1.0 10.5 15.07 139.9 0.70 0.48
4 14 2.0 09.0 11.17 151.2 0.72 0.51
4 14 4.0 08.6 09.84 153.6 0.72 0.51
4 14 6.0 08.8 09.46 151.1 0.71 0.52
5.4 Ablation Studies
We further dissect the effectiveness of ERW by conducting ablation studies on various design choices and parameter settings.
Target Representation.
We first compare alignment with multiple self-supervised encoders: MoCov3, CLIP, and DINOv2, as summarized in  Table 5.3.
  1. (a)
    Universality of Pretrained Encoders. All encoders tested offer improvements over baselines, indicating that ERW can broadly benefit from a variety of strong representation models.
  2. (b)
    Marginal Differences among DINOv2 Variants. DINOv2-B, DINOv2-L, and DINOv2-g yield comparable gains, suggesting that ERW does not require the largest possible teacher encoder for effective representation transfer.
Summary. Regardless of which high-quality self-supervised model is used, early-stage alignment provides a consistent boost in sample fidelity and training efficiency.
Placement of ERW Depth.
We hypothesize that early layers in the diffusion backbone primarily learn semantic features (the L2R circuit), whereas deeper layers specialize in generative decoding.
  1. (a)
    Empirical Validation. In  Table 5.3, initializing the earliest layers (0–8) notably outperforms re-initializing middle or late sections (FID 54.2 vs. $>67$).
  2. (b)
    Consistent with Circuit Perspective. This corroborates our three-stage diffusion circuit ( Section 3), underscoring that aligning deeper layers for representation can be suboptimal since those layers focus more on generative refinement.
Summary. Targeting the front portion of the network for warmup is crucial, reinforcing our theoretical claim that representation learning predominantly resides in early layers.
Projection Depth and Alignment Weight.
We also investigate how the final projection head depth and the alignment-loss coefficient $\lambda$ affect training ( Table 5.3).
  1. (a)
    Optimal Hyperparameters. Using 4 warmup layers, a projection head at depth 14, and $\lambda=4.0$ achieves an FID of 8.6 at 50k iterations—a substantial gain over baselines.
  2. (b)
    Trade-off in $\lambda$. Larger $\lambda$ offers stronger representation alignment initially but may disrupt convergence if pushed to extremes, highlighting the need for moderate scheduling.
Summary. Properly balancing projection depth and alignment strength is key to maximizing ERW ’s benefits in practice.
Representation Dynamics.
We examine the temporal progression of representation alignment in  Figure 5.
  1. (a)
    Initial Dip, Subsequent Recovery. Alignment falls early on as the pretrained features adjust to the diffusion objectives, but it then recovers and improves.
  2. (b)
    Role of Decaying Guidance. A decaying weight in the alignment term ( Section 4.2) fosters stable synergy between semantic alignment and generative refinement.
Summary. Representation alignment follows a U-shaped trajectory, revealing the model’s gradual reconfiguration of pretrained features for denoising tasks before distilling them into robust, generative-friendly embeddings.
CKNNA Analysis.
Finally, we measure layer-wise representation quality using Class-conditional $k$-Nearest Neighbor Accuracy (CKNNA) [3], which indicates how well the hidden features capture class discriminability.
  1. (a)
    Improved Semantic Alignment. Compared to random initialization, ERW yields systematically higher CKNNA scores, confirming stronger semantic preservation.
  2. (b)
    Evolving Layer-wise Semantics. The alignment initially drops then recovers, mirroring the trends seen in  Figure 5 and supporting the conclusion that pretrained features are effectively integrated rather than merely overwritten.
Summary. CKNNA measurements reinforce our findings that ERW preserves and refines semantic information across layers, enabling more discriminative feature representations to emerge during diffusion training.
(a) Alignment with ERW
(b) Training Dynamic for Alignment
Figure 5: Scalability of ERW. (a) Alignment with ERW plot showing that the representation model’s alignment with the ERW model exhibits a high degree of alignment across all ERW depths. (b) Training dynamics for alignment indicate that within the 500K training steps for SiT-XL/2, the alignment between DINOv2-g and the diffusion model first decreases and then increases.
5.5 Hyperparameter Tuning
We adopt a bisection-style search to determine the key hyperparameters for ERW, specifically the ERW Depth (i.e., which early layers to initialize), the Projection Depth, and the initial value of $\lambda$ in Eq. (14). To keep the search computationally manageable, we do the following for each candidate hyperparameter setting:
  • We run a short warmup stage for $10\mathrm{k}$ iterations, followed by $20\mathrm{k}$ iterations of main diffusion training.
  • To evaluate performance quickly, we reduce the sampling steps from the usual $250$ to $50$ and generate only $10\mathrm{k}$ samples (instead of $50\mathrm{k}$) to compute a preliminary FID score.
This procedure substantially reduces the search cost while retaining sufficient fidelity to guide hyperparameter choices. In practice, around three to five such tests suffice to converge upon near-optimal settings for ERW Depth, Projection Depth, and $\lambda$, enabling both efficient training and high-quality generation.
6 Conclusion and Future Work
In this work, we introduced Embedded Representation Initialization (ERW), a novel method that significantly enhances the training efficiency and representation quality of diffusion models by embedding high-quality representations from pretrained self-supervised encoders into the model’s representation processing region. Through both theoretical analysis and empirical validation, we demonstrated that ERW accelerates training convergence by up to 40$\times$ and improves the fidelity of generated data, establishing new state-of-the-art performance based on REPA [54]. Our findings underscore the importance of make representation embedded in initialization in optimizing diffusion model training.
Appendix A Loss Function Decomposition Proof
This section presents a theoretical decomposition of the diffusion model’s loss function into two orthogonal learning mechanisms: representation learning through latent-to-representation (L2R) mapping and generation inversion through representation-to-generation (R2G) dynamics. The decomposition reveals fundamental statistical independence between representation quality and generation fidelity in diffusion processes.
Let the diffusion model’s score matching objective be defined as:
$\mathcal{L}_{\text{diff}}=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|F_{\theta}(\mathbf{z}_{t},t)-\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})\|^{2}\right],$ (19)
where $F_{\theta}(\mathbf{z}_{t},t)$ parameterizes the estimated score function, and $\lambda(t)$ is a time-dependent weighting function.
Our analysis hinges on the existence of a sufficient representation $\mathbf{r}_{t}=R(\mathbf{z}_{t})$ that statistically decouples the latent space geometry from generation dynamics. Formally, we assume:
{assumptionframe}
Assumption 1 (Sufficient Representation) .
The representation $\mathbf{r}_{t}=R(\mathbf{z}_{t})$ is a sufficient statistic for $\mathbf{z}_{t}$ with respect to the data distribution $p(\mathbf{z}_{t})$, implying that $\mathbf{r}_{t}$ captures all necessary information about $\mathbf{z}_{t}$ for generation. This sufficiency leads to the decomposition of the Fisher score:
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})=\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})+\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t}),$ (20)
where the generation score $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})$ and the representation score $\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})$ are orthogonal under the Fisher-Rao metric:
$\mathbb{E}_{\mathbf{z}_{t}\mid\mathbf{r}_{t}}\left[\langle\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t}),\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})\rangle\right]=0.$ (21)
This orthogonality reflects the statistical independence between the generation mechanism and the representation learning process, ensuring that the two components can be optimized independently.
Under this assumption, the ground truth score decomposes into generation and representation components:
$\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t})=\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})}_{\text{Generation Score}}+\underbrace{\nabla_{\mathbf{z}_{t}}\log p(\mathbf{r}_{t})}_{\text{Representation Score}}.$ (22)
Parametrizing the model’s score function as $F_{\theta}(\mathbf{z}_{t},t)=\mathcal{G}_{\theta}(\mathbf{r}_{t})+\mathcal{R}_{\theta}(\mathbf{z}_{t})$, where $\mathcal{G}_{\theta}$ and $\mathcal{R}_{\theta}$ independently model generation and representation scores, we expand the loss function:
$\mathcal{L}_{\text{diff}}$ $=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|(\mathcal{G}_{\theta}(\mathbf{r}_{t})+\mathcal{R}_{\theta}(\mathbf{z}_{t}))-(\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})+\nabla\log p(\mathbf{r}_{t}))\|^{2}\right]$ (23)
$=\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{R}_{\theta}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t})\|^{2}\right]$
$\quad+\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{G}_{\theta}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\|^{2}\right]$
$\quad+2\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\langle\mathcal{R}_{\theta}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t}),\mathcal{G}_{\theta}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\rangle\right].$ (24)
The cross-term vanishes due to the orthogonality condition in Assumption 1, as:
$\mathbb{E}_{\mathbf{z}_{t}\mid\mathbf{r}_{t}}\left[\langle\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t}),\nabla\log p(\mathbf{r}_{t})\rangle\right]=0.$ (25)
This yields the decoupled learning objectives:
$\mathcal{L}_{\text{diff}}=\underbrace{\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{R}_{\theta}(\mathbf{z}_{t})-\nabla\log p(\mathbf{r}_{t})\|^{2}\right]}_{\mathcal{L}_{\text{L2R}}}+\underbrace{\mathbb{E}_{t,\mathbf{z}_{t}}\left[\lambda(t)\|\mathcal{G}_{\theta}(\mathbf{r}_{t})-\nabla\log p(\mathbf{z}_{t}\mid\mathbf{r}_{t})\|^{2}\right]}_{\mathcal{L}_{\text{R2G}}}.$ (26)
Through the above decomposition, we effectively split the optimization objective of the diffusion model into two independent parts: Representation Learning $\mathcal{R}_{\theta}(\mathbf{z}_{t})$ and Generation Inversion $\mathcal{G}_{\theta}(\mathbf{r}_{t})$.
Appendix B Analysis Details
B.1 Evaluation details
CKNNA (Centered Kernel Nearest-Neighbor Alignment) is a relaxed version of the popular Centered Kernel Alignment (CKA; Kornblith et al. 25) that mitigates the strict definition of alignment. We generally follow the notations in the original paper for an explanation [19].
First, CKA have measured global similarities of the models by considering all possible data pairs:
$\mathrm{CKA}(\mathbf{K},\mathbf{L})=\frac{\mathrm{HSIC}(\mathbf{K},\mathbf{L})}{\sqrt{\mathrm{HSIC}(\mathbf{K},\mathbf{K})\mathrm{HSIC}(\mathbf{L},\mathbf{L})}},$ (27)
where $\mathbf{K}$ and $\mathbf{L}$ are two kernel matrices computed from the dataset using two different networks. Specifically, it is defined as $\mathbf{K}_{ij}=\kappa(\phi_{i},\phi_{j})$ and $\mathbf{L}_{ij}=\kappa(\psi_{i},\psi_{j})$ where $\phi_{i},\phi_{j}$ and $\psi_{i},\psi_{j}$ are representations computed from each network at the corresponding data $\mathbf{x}_{i},\mathbf{x}_{j}$ (respectively). By letting $\kappa$ as a inner product kernel, $\mathrm{HSIC}$ is defined as
$\mathrm{HSIC}(\mathbf{K},\mathbf{L})=\frac{1}{(n-1)^{2}}\Big{(}\sum_{i}\sum_{j}\big{(}\langle\phi_{i},\phi_{j}\rangle-\mathbb{E}_{l}[\langle\phi_{i},\phi_{l}\rangle]\big{)}\big{(}\langle\psi_{i},\psi_{j}\rangle-\mathbb{E}_{l}[\langle\psi_{i},\psi_{l}\rangle]\big{)}\Big{)}.$ (28)
CKNNA considers a relaxed version of Eq. (27) by replacing $\mathrm{HSIC}(\mathbf{K},\mathbf{L})$ into $\mathrm{Align}(\mathbf{K},\mathbf{L})$, where $\mathrm{Align}(\mathbf{K},\mathbf{L})$ computes Eq. (28) only using a $k$-nearest neighborhood embedding in the datasets:
$\mathrm{Align}(\mathbf{K},\mathbf{L})=\frac{1}{(n-1)^{2}}\Big{(}\sum_{i}\sum_{j}\alpha(i,j)\big{(}\langle\phi_{i},\phi_{j}\rangle-\mathbb{E}_{l}[\langle\phi_{i},\phi_{l}\rangle]\big{)}\big{(}\langle\psi_{i},\psi_{j}\rangle-\mathbb{E}_{l}[\langle\psi_{i},\psi_{l}\rangle]\big{)}\Big{)},$ (29)
where $\alpha(i,j)$ is defined as
$\alpha(i,j;k)=\mathbb{1}[i\neq j\,\,\text{and}\,\,\phi_{j}\in\mathrm{knn}(\phi_{i};k)\,\,\text{and}\,\,\psi_{j}\in\mathrm{knn}(\psi_{i};k)],$ (30)
so this term only considers $k$-nearest neighbors at each $i$. In this paper, we randomly sample 10,000 images in the validation set in ImageNet [7] and report CKNNA with $k=10$ based on observation in Huh et al. [19] that smaller $k$ shows better a better alignment.
B.2 Description of pretrained visual encoders
  • MoCov3 [6] studies empirical study to train MoCo [13, 5] on vision transformer and how they can be scaled up.
  • CLIP [42] proposes a contrastive learning scheme on large image-text pairs.
  • DINOv2 [39] proposes a self-supervised learning method that combines pixel-level and patch-level discriminative objectives by leveraging advanced self-supervised techniques and a large pre-training dataset.
Appendix C Hyperparameter and More Implementation Details
Table 6: Hyperparameter setup.
Figure 1,2,3 Table 3,4 (SiT-B) Table 1,2,5 (SiT-XL)
Architecture
Input dim. 32$\times$32$\times$4 32$\times$32$\times$4 32$\times$32$\times$4
Num. layers 28 12 24
Hidden dim. 1,152 768 1,152
Num. heads 16 12 16
ERW
$\mathrm{sim}(\cdot,\cdot)$ NT-Xent NT-Xent NT-Xent
Encoder $f(\mathbf{x})$ DINOv2-B DINOv2-B DINOv2-B
Optimization
Batch size 256 256 256
Optimizer AdamW AdamW AdamW
lr 0.0001 0.0001 0.0001
$(\beta_{1},\beta_{2})$ (0.9, 0.999) (0.9, 0.999) (0.9, 0.999)
Interpolants
$\alpha_{t}$ $1-t$ $1-t$ $1-t$
$\sigma_{t}$ $t$ $t$ $t$
$w_{t}$ $\sigma_{t}$ $\sigma_{t}$ $\sigma_{t}$
Training objective v-prediction v-prediction v-prediction
Sampler Euler-Maruyama Euler-Maruyama Euler-Maruyama
Sampling steps 250 250 250
Guidance - - -
Further implementation details. We implement our model based on the original SiT implementation [36]. Throughout the experiments, we use the exact same structure as DiT [40] and SiT [36]. We use AdamW [24, 35] with constant learning rate of 1e-4, $(\beta_{1},\beta_{2})=(0.9,0.999)$ without weight decay. To speed up training, we use mixed-precision (fp16) with a gradient clipping. We also pre-compute compressed latent vectors from raw pixels via stable diffusion VAE [44] and use these latent vectors. Because of this, we do not apply any data augmentation, but we find this does not lead to a big difference, as similarly observed in EDM2 [22]. We also use stabilityai/sd-vae-ft-ema decoder for decoding latent vectors to images. For MLP used for a projection, we use three-layer MLP with SiLU activations [9]. We provide a detailed hyperparameter setup in   Table 6.
Pretrained encoders. For MoCov3-B and -L models, we use the checkpoint in the implementation of RCG [32]; [^1] https://github.com/LTH14/rcg for other checkpoints, we use their official checkpoints released in their official implementations. To adjust a different number of patches between the diffusion transformer and the pretrained encoder, we interpolate positional embeddings of pretrained encoders.
Sampler. For sampling, we use the Euler-Maruyama sampler with the SDE with a diffusion coefficient $w_{t}=\sigma_{t}$. We use the last step of the SDE sampler as 0.04, and it gives a significant improvement, similar to the original SiT paper [36].
Training Tricks. We explore the influence of various training techniques on ERW’s performance. Notably, we observe performance improvements when incorporating Rotary Positional Embeddings [50].
Table 7: Impact of Training Tricks in ERW. Using the SD-VAE [44], ERW achieves an FID of 55.6 at 50K training steps on ImageNet class-conditional generation. This table illustrates how each training trick incrementally improves the FID, demonstrating that advanced design techniques enhance the original DiT performance.
Training Trick Training Step FID-50$\downarrow$
Representation Alignment Loss
+ REPA [54] 50K 78.2
Architecture Improvements
+ Rotary Pos Embed [50] 50K 73.6
Initialization
+ ERW (Ours) 50K 51.7
Discussion of Quantifiable Metrics.
In addition to the cost breakdown above, the following metrics are instrumental for rigorous quantitative analysis:
  • Iteration Time ($t_{\mathrm{iter}}$): The average time per training iteration (in seconds), which directly impacts $C_{w}$.
  • Warmup Iteration Count ($N_{w}$): The precise number of iterations employed during the warmup phase.
  • Main Training Iteration Count ($N_{m}$): The total number of iterations in the main training phase.
  • Hyperparameter Search Steps ($b_{i}$): The number of binary search steps needed for each hyperparameter.
  • Evaluation Efficiency: The use of a reduced evaluation protocol (e.g., FID measurement on 10K images using 50 sampling steps) accelerates the hyperparameter search process.
  • Overall Speedup Factor: The ratio $C_{m}/C_{w}$ is approximately 10, underscoring that the main training phase is the dominant cost.
Appendix D Evaluation Details
We strictly follow the setup and use the same reference batches of ADM [8] for evaluation, following their official implementation. [^2] https://github.com/openai/guided-diffusion/tree/main/evaluations We use 8$\times$×NVIDIA H800 80GB GPUs or for evaluation and enable tf32 precision for faster generation, and we find the performance difference is negligible to the original fp32 precision.
In what follows, we explain the main concept of metrics that we used for the evaluation.
  • FID [15] measures the feature distance between the distributions of real and generated images. It uses the Inception-v3 network [51] and computes distance based on an assumption that both feature distributions are multivariate gaussian distributions.
  • sFID [37] proposes to compute FID with intermediate spatial features of the Inception-v3 network to capture the generated images’ spatial distribution.
  • IS [46] also uses the Inception-v3 network but use logit for evaluation of the metric. Specifically, it measures a KL-divergence between the original label distribution and the distribution of logits after the softmax normalization.
  • Precision and recall [27] are based on their classic definitions: the fraction of realistic images and the fraction of training data manifold covered by generated data.
Appendix E Baselines
In what follows, we explain the main idea of baseline methods that we used for the evaluation.
  • ADM [8] improves U-Net-based architectures for diffusion models and proposes classifier-guided sampling to balance the quality and diversity tradeoff.
  • VDM++ [23] proposes a simple adaptive noise schedule for diffusion models to improve training efficiency.
  • Simple diffusion [18] proposes a diffusion model for high-resolution image generation by exploring various techniques to simplify a noise schedule and architectures.
  • CDM [17] introduces cascaded diffusion models: similar to progressiveGAN [21], it trains multiple diffusion models starting from the lowest resolution and applying one or more super-resolution diffusion models for generating high-fidelity images.
  • LDM [44] proposes latent diffusion models by modeling image distribution in a compressed latent space to improve the training efficiency without sacrificing the generation performance.
  • U-ViT [2] proposes a ViT-based latent diffusion model that incorporates U-Net-like long skip connections.
  • DiffiT [12] proposes a time-dependent multi-head self-attention mechanism for enhancing the efficiency of transformer-based image diffusion models.
  • MDTv2 [11] proposes an asymmetric encoder-decoder scheme for efficient training of a diffusion-based transformer. They also apply U-Net-like long-shortcuts in the encoder and dense input-shortcuts in the decoder.
  • MaskDiT [55] proposes an asymmetric encoder-decoder scheme for efficient training of diffusion transformers, where they train the model with an auxiliary mask reconstruction task similar to MAE [14].
  • SD-DiT [56] extends MaskdiT architecture but incorporates self-supervised discrimination objective using a momentum encoder.
  • DiT [40] proposes a pure transformer backbone for training diffusion models based on proposing AdaIN-zero modules.
  • SiT [36] extensively analyzes how DiT training can be efficient by moving from discrete diffusion to continuous flow-based modeling.